solve each absolute value equation

6. |3x+2|+3=0
1
Expert's answer
2012-12-10T06:01:06-0500
solve each absolute value equation
&
6. |3x+2|+3=0

|3x+2|+3 = 0& ==>

|3x+2| = -3& ==>

(3x+2) є {empty set},

cause |·| is a non-negative function by the definition. So, the given equation has no possible solutions.